11. Find 8th term of the sequence 1,5,9.......

12. If central angle of an arc is 200º,then find the angle made by the arc on the

alternate arc?

13. Prove that a rectangle is always cyclic

14. If algebraic form of an arithmetic sequence is 3-5n,find first term and common

difference

15. Find number of multiples of 11 in between 100 a nd 200


1
Expert's answer
2022-01-03T17:10:54-0500

11. It is the arithmetic sequence, so an=a1+d*(n-1).

a1=1, d=a2-a1=5-1=4.

a8=a1+7*d=1+7*4=29.

Answer: 8th term of the sequence is 29.


12. The angle made by an arc of a circle on

the alternate arc is half the angle made at

the centre. So if central angle is 200°, then answer will be 200°*1/2=100°.


13. Any rectangle have all it's angles 90° each. The quadrilateral is cyclic, if the opposite angles that lie on the same diagonal add 180°. In any rectangle sum of opposite angles will be 90°+90°=180°. So any rectangle is cyclic quadrilateral.


14. First term: a1=3-5*1=-2;

Second term: a2=3-5*2=-7;

Common difference: d=a2-a1=-7-(-2)=-5.

Answer: first term is -2 and common difference is -5.


15. 100 < 11n < 200;

100/11 < n < 200/11;

9 1/11 < n < 18 2/11;

n=10.. 18 (n is a whole number from 10 to 18). From 10 to 18 is total 9 whole numbers.

Answer: 9 multiples of 11 is in between 100 and 200.